OpenStudy (anonymous):

A ladder 13ft long rests against a vertical wall and is sliding down the wall at a rate of 3ft/s at the instant the foot of the ladder is 5 feet from the base of the wall. At this instant, how fast is the foot of the ladder moving away from the wall?

OpenStudy (anonymous):

thanks for your help

OpenStudy (bahrom7893):

 Sorry I kept crashing

OpenStudy (anonymous):

yeah i got that, and then y=12feet at the instant it is 5 feet from the base

OpenStudy (bahrom7893):

\[z^2=x^2+y^2\] \[2z\frac{dz}{dt}=2x\frac{dx}{dt}+2y\frac{dy}{dt}\] \[z\frac{dz}{dt}=x\frac{dx}{dt}+y\frac{dy}{dt}\]

OpenStudy (bahrom7893):

\[\frac{dz}{dt}=0\] as the length of the ladder doesn't change, so: \[0=x*\frac{dx}{dt}+y*\frac{dy}{dt}\]

OpenStudy (anonymous):

yeah i follow, and since z^2 is 144, and its asking for dy/dt then we evaluate that equation to get \[\frac{ dy }{ dt } = -\frac{ x }{ y }(\frac{ dx }{ dt })\]

OpenStudy (bahrom7893):

Yea, just use the Pythagorean theorem to find y, and you know that \[\frac{dy}{dt}=-3\frac{ft}{s}\] since it's sliding down

OpenStudy (bahrom7893):

x is given

OpenStudy (anonymous):

\[-3(\frac{ -12 }{ 5 }) = \frac{ dx }{ dt }\]

OpenStudy (anonymous):

correct?

OpenStudy (bahrom7893):

Can you post the intermediate steps? hahah I'm too lazy to evaluate.

OpenStudy (anonymous):

\[\frac{ dx }{ dt } = \frac{ -5 }{ 4 }\]

OpenStudy (anonymous):

y=12, because its a 13,12,5 triangle

OpenStudy (bahrom7893):

looks correct to me

OpenStudy (anonymous):

alright, just to double check the origional equation is asking for dx/dt?

OpenStudy (bahrom7893):

How fast is the foot of the wall, moving away, so how fast dx/dt is increasing.. Wait it should have been positive, why did you get negative?

OpenStudy (dan815):